OpenStudy (anonymous):

Prove that n^2+5n+2 is an even integer for all n<-N

OpenStudy (freckles):

what does n<-N mean?

OpenStudy (freckles):

if you meant for n in the natural numbers then you can do this by induction

OpenStudy (freckles):

if n=1 then n^2+5n+2=? <--is this even Now assume for some integer k>=1 k^2+5k+2 can be expressed as 2i we want to show (k+1)^2+5(k+1)+2 can also be expressed as 2l l and i are integers by the way

OpenStudy (freckles):

you can also do this without induction

OpenStudy (freckles):

if you know 2|n(n+1)

OpenStudy (freckles):

\[n^2+5n+2 \\ n^2+n+4n+2 \\ n(n+1)+4n+2\] so if you know 2|n(n+1) then you can play with that way and it isn't so bad :)

OpenStudy (anonymous):

this is what i meant \[n^2+5n+2, n \ \leftarrow \mathbb{N}\]

OpenStudy (anonymous):

If i use induction, do i do two cases, even and odd or jut even? @freckles

OpenStudy (freckles):

we are trying to show it is even so we are trying to show n^2+5n+2 is even for any integer n and so we only need to show it is just for n=1 then assume it is true for n=k and show it is true for n=k+1 so yes it is a one case type of proof

OpenStudy (freckles):

\[(k+1)^2+5(k+1)+2\] hint here do a little multiplying so you use k^2+5k+2=2i

OpenStudy (anonymous):

okay i got it. thanks

OpenStudy (freckles):

np